A ball is thrown




horizontally with an initial velocity of




30m/s from a height of 2 m. How long will it take to reach the ground? How far will it go?

1
Expert's answer
2022-04-18T08:50:12-0400

The time of motion

"t=\sqrt{\frac{2H}{g}}=\sqrt{\frac{2*2}{9.8}}=0.64\:\rm s"

The horizontal distance

"d=v_0t=30*0.64=19\:\rm m"